# Don’t be Fooled: The Misinformation Effect of Explanations in Human-AI Collaboration

### v1

@@ 3. Theoretical Development
In the evolving field of I, understanding the impact of AI explanations on decision-making has become critical. Explanations serve as a bridge between AI and humans, influencing trust, reliance, and collaboration (Schemmer et al., 2023c; Hemmer et al., 2023; Schoeffer et al., 2024). This work investigates how incorrect explanations, when paired with accurate AI advice, can mislead humans, potentially impairing their procedural knowledge and reasoning capabilities.
Schemmer et al. (2023c) highlight the role of AI advice in AI-assisted decision-making, emphasizing how they shape humans’ reliance on AI. Next to the advice, explanations influence the collaboration between humans and AI. Morrison et al. (2024) extend this understanding, exploring the effects of both correct and incorrect explanations on human reliance on AI. Their findings suggest that even when AI advice is accurate, incorrect explanations can distort human perceptions and decision-making processes, leading to an inappropriate reliance on AI. Building on these insights, our research focuses on a specific aspect of the work of Morrison et al. (2024): the misleading potential of incorrect explanations in the context of correct AI advice. While prior research has taken the first steps towards investigating the impact of incorrect explanations, still little is known about the repercussions on the ability of human recipients to perform tasks, namely their procedural knowledge, and to conclude based on their understanding, namely their reasoning, after the collaboration with the AI. We examine how such flawed explanations affect humans’ procedural knowledge, reasoning, and overall human-AI collaboration, contributing new insights to the growing body of literature on AI-assisted decision-making.
In AI-assisted decision-making, declarative knowledge (the ”know-what”) and procedural knowledge (the ”know-how”) are both crucial. Declarative knowledge is factual knowledge or information that humans hold. When exposed to incorrect explanations, incorrect explanations can distort this knowledge base, leading to flawed reasoning, as posited by Stenning and Van Lambalgen (2012).
Procedural knowledge refers to the ability to perform tasks and make decisions. The acquisition of procedural knowledge is highly sensitive to cognitive load, as discussed by Sweller (1988). Incorrect explanations can not only impair procedural knowledge, but also increase cognitive load, leading to difficulties in learning and applying procedures effectively. Chi and Wylie (2014) further support this by highlighting the importance of cognitive engagement in learning, which can be hindered by misinformation. Given these dynamics, we assume that incorrect explanations impair humans’ procedural knowledge. This is supported by Roediger and Butler (2011), who highlight how incorrect information encountered during learning can lead to the retention of false information, thereby impairing reasoning and memory. Similarly, Johnson and Seifert (1994) argue that incorrect information, even after correction, can still lead to wrong inferences. With prior research emphasizing the negative repercussions on procedural knowledge, we hypothesize:
@@
Incorrect explanations lead to lower procedural knowledge.
When collaborating with AI, incorrect explanations may impair humans’ reasoning capabilities as they transition from knowing ”how” to understanding ”why”. In psychology, the work of Loftus et al. (1978) shows that misleading information can impair human memory, leading to an erroneous understanding. Ecker et al. (2011) demonstrate the persistence of misinformation effects, showing how they continue to influence reasoning even after correction, thus highlighting their lasting impact. Similarly, Kendeou et al. (2013) investigate the effects of incorrect information on human reasoning and explore mechanisms to mitigate these effects. Building upon prior research that examines how misinformation impairs decision-making and reasoning (Soon and Goh, 2018), we adopt this perspective to AI-assisted decision-making and hypothesize the following:
@@
The reasoning for the classification in the task is impaired by incorrect explanations.
Seeber et al. (2020) emphasize the importance of communication for effective collaboration between humans and AI. They argue that misunderstandings or unclear communication can lead to disruptions in team performance, particularly when AI systems provide incorrect or unclear explanations. Furthermore, Dzindolet et al. (2003) investigate how trust in automated systems is impacted by incorrect feedback, leading to reduced team performance. Similarly, Eiband et al. (2018) provide insights into how human-AI collaboration is affected by the transparency of AI systems, suggesting that incorrect advice can hinder effective teamwork. Lastly, Bansal et al. (2021) explore the dynamics of human-AI interaction, showing that incorrect AI advice alongside explanations can create friction in collaboration, leading to poorer outcomes. Incorrect explanations in AI-assisted decision-making can impair the collaboration, leading to inappropriate reliance on AI (Morrison et al., 2024). With prior research showing the relationship between humans’ reliance behavior on AI and the human-AI team performance (Schemmer et al., 2023c), we assume that the human-AI team performance drops when humans are provided with incorrect explanations:
@@
The collaboration with a misleading AI leads to a lower human-AI team performance.
@@ 4. Methodology
In this section, we describe our methodology to assess how incorrect explanations for correct AI advice influence humans during and post collaboration with an AI. We set up an online study and investigated how participants performed in a visual classification task on an architectural dataset. In this section, we outline the task domain, the study design, the recruitment of participants, the development of the AI, and finally, the metrics that we use to assess our RQs. Before we ran the study, we pre-registered the study on AsPredicted.org to report our hypotheses, our treatments, our planned analyses, and our exclusion strategy. An anonymized copy of the pre-registration is provided in the supplemental materials.
@@ 4.1. Task Domain
In order to analyze the impact of explanations, we chose to use a task that most people are not familiar with and typically cannot handle themselves initially: the classification of the architectural style of buildings. To do so, we use the established dataset of Xu et al. (2014) containing images of buildings across 25 different architectural styles. In close discussion with architecture researchers of the local university, we chose three architectural styles that share similar features and are not easily distinguishable: Art Nouveau, Art Deco, and Georgian Architecture. For each architectural style, we selected 30 images that clearly represent the features of each architectural style and, thus, are appropriate instances for our study. We further made sure that the buildings were centered in each image and cropped all irrelevant information in the images, like other buildings.
@@ 4.2. Study Design
The online study was divided into five different parts (see Figure 1): i part (1), the participants had to give their consent to participate and were introduced to the study and its procedure. They also received context information about the three different architectural styles with a description of their main characteristics (see Table 3 in Appendix A). Additionally, we included two attention checks, one of them in part (1): we asked participants what their task would be throughout the study, and they could select from three options. In part (2), a pre-test assessed participants’ task performance in classifying architectural styles of buildings: they were each shown six images randomly drawn from a bucket of 18 pre-selected images of the dataset. We ensured that each class was balanced for each participant (two images per class) and that the order in which the classes were shown to participants was varied. By doing so, we minimized the risk of inducing biases in the order of images. It also ensured that our results were not dependent on the difficulty of pre-selected images. Moreover, the pre-test allowed us to check that participants in each treatment did not differ in their prior expertise regarding the task. Following the pre-test, participants were randomly assigned to a treatment in the main task in part (3), distinguishing the type of AI support received:
@@
To prevent participants from saving time by just randomly clicking through the study, participants could continue to the next page in the pre-test, main task, and post-test only after a few seconds. This design choice followed the protocol of (Spitzer et al., 2024a) and was to ensure that participants focused on the actual task. In the final part (5), participants answered a questionnaire. The questionnaire asked them to describe and explain each architectural style. With this, we were able to analyze whether participants had learned the correct distinctions between the architectural styles and were able to conclude based on their understanding (e.g., their reasoning). On top of that, they had to rate several variables on seven-point Likert scales to measure for confounding factors. These included the AI usefulness, their experience with AI, their cognitive load, and their trust in AI. All variables are presented in Appendix A in Table 5. Throughout the questionnaire, we implemented a second attention check to ensure only valid results, as suggested by Abbey and Meloy (2017). In this attention check, we asked participants to select ”Likey“ for an item of AI usability (”As an attention check, please choose ”Likely” for this statement“). The questionnaire ended with demographic questions on participants’ age, gender, education and employment.
@@ 4.3. AI Development
The AI used in the study was a large language model (LLM) that provided participants with a classification (AI advice) and an explanation for its reasoning, depending on the treatment they were assigned to. As LLM, we used Open AI’s GPT-4o model (model version 2024-05-13) through an Azure Open AI Studio instance. In the pre-phase of the study, we tested multiple prompt strategies in a workshop with three authors. The final prompts that were used are shown in Appendix A in Table 4. The LLM was prompted such that the correct explanation treatment provided an explanation that corresponded to its prediction, while the explanation did not match the prediction in the incorrect explanation treatment.
The definition of an incorrect explanation follows Morrison et al. (2024) in that the AI does not provide correct reasoning for its predicted class, independent of the ground truth of the image. For instance, if the AI made a correct prediction for an Art Nouveau building, but the explanation does not conform to this class, it is defined as incorrect. To avoid information overload (Arnold et al., 2023), we designed the explanations to be no longer than three sentences. As outlined in Section 3, we analyze the impact of incorrect explanations for cases where the AI classification is correct.
@@ 4.5. Metrics
Similar to previous work (e.g., (Hemmer et al., 2023; Schoeffer et al., 2024)), we assessed participants’ task performance in classifying the architectural styles in pre-test, main task, and post-test. The task performance was used to approximate participants’ procedural knowledge in the pre-test and post-test and their human-AI team performance in the main task. As metrics for the task performance, we used accuracy and measured the ratio of correctly classified images over all images. Participants had to select one of the three different architectural styles for each image by selecting from a drop-down menu on each page of the task. This means that a random guess corresponded to 33.3% of performance. Aligning with Schoeffer et al. (2024), we also measured the correct adherence and detrimental overrides in the main task of the study.
We assessed participants’ reasoning ability for the three architectural styles through open-ended questions by asking them to describe and explain each style, thereby following the procedure of Chi et al. (1989). We rated each answer in terms of correctness by comparing it to the correct characteristics and features for each architectural style. By doing so, we were able to assess whether incorrect explanations impaired participants’ reasoning.
Finally, we established several control variables to investigate the potential underlying factors that might influence AI-assisted decision-making. In particular, we controlled for participants’ cognitive load as previous research suggests that the information in explanations displayed to humans can affect their decision-making behavior (Abdul et al., 2020; Hudon et al., 2021; Herm, 2023; Spitzer et al., 2024a). We measured participants’ cognitive load on a seven-point Likert scale by having them rate five validated items previous research has established (Senoner et al., 2024). In addition, we assessed participants’ AI trust, AI usefulness, and experience with AI by using items proven in previous research (Senoner et al., 2024), all of them also rated on a seven-point Likert scale. All items are shown in Table 5 in Appendix A.
@@ 5. Results
To address the research questions, we first conduct several statistical analyses to answer RQ: 1 in subsection 5.1. Subsequently, we address RQ: 2 and qualitatively assess the open-ended questionnaires in subsection 5.2. In the final subsection 5.3, we evaluate the impact of incorrect explanations on the human-AI team performance (RQ: 3).
@@ 5.1. RQ1: Impact of Incorrect Explanations on Procedural Knowledge
It took participants, on average, 14 minutes and 54 seconds to complete the study. Overall, 160 participants passed the attention check and finished the study according to the study protocol. Of these 160 participants, 79 were male, 77 were female, and four identified as diverse. To establish a baseline and ensure that all treatments began on an equal performance level in classifying the architectural styles, we conduct a one-way ANOVA on the pre-test performance scores across the four treatments (). These results fail to reject the null hypothesis, indicating no significant differences in pre-test performance among the four treatments. This finding suggests that participants across all treatments start with comparable levels of procedural knowledge to classify the architectural styles prior to the main task. This allows for a thorough interpretation of any differences observed in the post-test results, as they can be more readily attributed to the treatment effects rather than pre-existing differences.
@@
To assess the significance of differences in post-test performance across treatments, we compare performance levels between treatments. A one-way ANOVA yields evidence of treatment effects (), indicating that the type of AI support in the main task influences participant performance in the post-test. To further examine these differences, we conduct pairwise comparisons using one-sided t-tests, assuming participants in the incorrect explanation treatment exhibit lower performance compared to participants in the other performance. We also report the corrected p-values according to the Holm-Bonferroni correction (Holm, 1979) (see Table 8 in Appendix A). There is no significant difference between the post-test performances in the incorrect explanation treatment and the control group (). Comparison with the AI treatment also shows no significant difference (). Notably, the comparison between incorrect and correct explanations yields a significant difference (), with incorrect explanations leading to inferior performance. With these findings, we find support for hypothesis 1. These results demonstrate that the accuracy of explanations plays a crucial role in AI-assisted decision-making, with incorrect explanations decreasing the task performance of humans in the post-test the most. Correct AI-generated explanations provide an additional performance advantage over incorrect explanations, underscoring the importance of explanation accuracy in AI-assisted tasks.
@@
To account for potential confounding factors influencing the effect of the different AI supports on procedural knowledge, we run regression analyses with the post-test performance as dependent variables and model the type of AI support as the independent variable. We define cognitive load, AI trust, AI knowledge, and AI usefulness as confounding factors. The regression analysis examining the impact of different treatments on post-test performance reveals that the type of AI support significantly influences participants’ performance (see Table 9 in Appendix A). The model accounts for of the variance in post-test performance (). Specifically, participants in the correct explanation treatment exhibit the greatest improvement in post-test performance, with a significant positive effect (), followed by those in the AI treatment () and the incorrect explanation treatment (). The data supports the findings of the statistical tests. Interestingly, cognitive load is negatively associated with post-test performance (), suggesting that higher cognitive load impedes procedural knowledge.
@@
Additionally, we analyze how the procedural knowledge develops from pre-test to post-test. We plot this difference in Figure 4. The figure illustrates whether procedural knowledge develops or degrades across different decision-making stages in the different treatments. Interestingly, the data reveals that the incorrect explanation treatment still leads to a procedural knowledge development. Conversely, the AI treatment and correct explanation treatment show the highest procedural knowledge development, exceeding the development in the incorrect explanation treatment.
@@
Furthermore, we run an ANOVA and Tukey HSD Post-Hoc Test which reveal significant differences in procedural knowledge gains across treatments (). Notably, the incorrect explanation treatment does not significantly differ from the AI treatment or the correct explanation treatment ( for both comparisons), but it shows a trend toward higher procedural knowledge gains compared to the control group (). The procedural knowledge development is significantly higher than in the control group (). This suggests that while incorrect explanations may increase the procedural knowledge in collaboration with an AI, the correctness of explanations reduces the effect compared to correct explanations.

### v2

@@ 3. Theoretical Development
In the evolving field of HCI, understanding the impact of AI explanations on decision-making has become critical. Explanations can serve as a bridge between AI and humans, influencing trust, reliance, and collaboration (Schemmer et al., 2023b; Hemmer et al., 2023; Schoeffer et al., 2024). This work investigates how incorrect explanations, when paired with accurate AI advice, can mislead humans, potentially impairing their procedural knowledge and reasoning capabilities. In Section 2, we present several works that investigate the impact of incorrect AI advice and incorrect explanations in decision-making. Building on prior research (Morrison et al., 2024; Cabitza et al., 2024a; Papenmeier et al., 2019), we derive several hypotheses grounded in related psychology and human behavior research to study the impact on humans’ knowledge.
Research in behavioral science distinguishes between declarative knowledge (the “know-what”) and procedural knowledge (the “know-how”) in decision-making (Herz and Schultz Jr, 1999). While both types of knowledge are interconnected, our study focuses on procedural knowledge as we want to determine the impact of incorrect explanations on humans’ downstream task performance. Humans’ ability to complete tasks effectively remains paramount, especially where AI support is not perfectly reliable or may not always be available (Laux, 2023). By measuring task performance, we can capture procedural knowledge (McCormick, 1997; Nahdi and Jatisunda, 2020) and the extent to which humans effectively apply their understanding in practice (Clark and Dumas, 2016). Exposure to incorrect explanations can significantly distort both declarative and procedural knowledge. As procedural knowledge is particularly vulnerable to cognitive disruptions (Sweller, 1988), incorrect explanations might increase cognitive load and impair the acquisition of procedural knowledge. Chi and Wylie (2014) provide crucial insights into this phenomenon, demonstrating how cognitive engagement can be systematically undermined by misleading information. Thus, we assume that incorrect explanations for correct AI advice impair humans’ procedural knowledge. We hypothesize:
@@
Incorrect explanations for correct AI advice lead to lower procedural knowledge.
While our study primarily centers on procedural knowledge, we also assess participants’ reasoning abilities to explore how individuals articulate the cognitive principles underlying their decision-making processes (Johnson and Seifert, 1994). This enables us to examine the extent to which incorrect explanations influence both task performance and conceptual understanding. Prior research has demonstrated the profound impact of misleading information on cognitive processes, revealing significant impairments in memory and comprehension (Loftus et al., 1978; Soon and Goh, 2018). Critically, these cognitive distortions persist even after subsequent correction attempts (Ecker et al., 2011; Kendeou et al., 2013) demonstrating the effect of misinformation on cognition and their potential to impair reasoning (Ecker et al., 2011). In the context of AI-assisted classification tasks, these prior findings suggest that incorrect explanations are likely to disrupt reasoning ability in classification tasks.
@@
Human reasoning is impaired by incorrect explanations for correct AI advice.
Seeber et al. (2020) emphasize the importance of communication between humans and AI for effective collaboration. They argue that misunderstandings or unclear communication can lead to disruptions in team performance, particularly when AI systems provide incorrect or unclear explanations. Furthermore, Dzindolet et al. (2003) investigate how trust in automated systems is impacted by incorrect feedback, leading to reduced team performance. Similarly, Eiband et al. (2018) provide insights into how human-AI collaboration is affected by the transparency of AI systems, suggesting that incorrect advice can hinder effective teamwork. Lastly, Bansal et al. (2021) explore the dynamics of human-AI interaction, showing that incorrect AI advice alongside explanations can create friction in collaboration, leading to poorer outcomes. Incorrect explanations in AI-assisted decision-making can impair the collaboration, leading to inappropriate reliance on AI (Morrison et al., 2024). With prior research showing the relationship between humans’ reliance behavior on AI and the human-AI team performance (Schemmer et al., 2023b), we assume that the human-AI team performance drops when humans are provided with incorrect explanations. Thus, we extend prior research by directly assessing the impact of incorrect explanations for correct AI advice and hypothesize:
@@
Incorrect explanations for correct AI advice lead to a lower human-AI team performance.
@@ 4. Methodology
In this section, we describe our methodology to assess how incorrect explanations for correct AI advice influence humans during and post collaboration with an AI. We set up an online study and investigated how participants performed in a visual classification task on an architectural dataset. We outline the task domain, the study design, the recruitment of participants, the development of the AI, and finally, the metrics that we use to assess our RQs. Before we ran the study, we pre-registered the study on AsPredicted.org to report our hypotheses, our treatments, our planned analyses, and our exclusion strategy. An anonymized copy of the pre-registration is provided in the supplemental materials.
@@ 4.1. Task Domain
To analyze the impact of explanations, we chose a task where most people have little experience, i.e., have no expert-level knowledge and typically cannot perform well themselves initially: the classification of the architectural style of buildings. This task represents various real-world scenarios in which AI is utilized: seeking advice and explanations for a task that is either unknown or humans might not possess enough knowledge to solve the task confidently. Such scenarios in real-world use cases range from providing medical diagnoses for patients to generating code for people with no background in programming. Thus, the study of how humans integrate incorrect explanations into their decision-making likely extends to more complex generative AI interactions. We use the established dataset of Xu et al. (2014) containing images of buildings across 25 different architectural styles. In close discussion with architecture researchers of the local university, we chose three architectural styles that share similar features and are not easily distinguishable: Art Nouveau, Art Deco, and Georgian Architecture. For each architectural style, we selected 30 images that clearly represent the features of each architectural style and, thus, are appropriate instances for our study. We further made sure that the buildings were centered in each image and cropped all irrelevant information in the images, like other buildings.
@@ 4.2. Study Design
The online study was divided into five different parts (see Figure 1): in part (1), the participants had to give their consent to participate and were introduced to the study and its procedure. They also received context information about the three different architectural styles with a description of their main characteristics (see Table 3 in Appendix A). Additionally, we included two attention checks, one of them in part (1): we asked participants what their task would be throughout the study, and they could select from three options. In part (2), a pre-test assessed participants’ task performance in classifying architectural styles of buildings: they were each shown six images randomly drawn from a bucket of 18 pre-selected images of the dataset. We ensured that each class was balanced for each participant (two images per class) and that the order in which the classes were shown to participants was varied. By doing so, we minimized the risk of inducing biases in the order of images. It also ensured that our results were not dependent on the difficulty of pre-selected images. Moreover, the pre-test allowed us to check that participants in each treatment did not differ in their prior expertise regarding the task. Following the pre-test, participants were randomly assigned to a treatment in the main task in part (3), distinguishing the type of AI support received:
@@
To prevent participants from saving time by just randomly clicking through the study, participants could continue to the next page in the pre-test, main task, and post-test only after a few seconds. This design choice followed the protocol of (Spitzer et al., 2024a) and was to ensure that participants focused on the actual task. In the final part (5), participants answered a questionnaire. The questionnaire asked them to describe and explain each architectural style. With this, we were able to analyze whether participants had learned the correct distinctions between the architectural styles and were able to conclude based on their understanding (e.g., their reasoning). On top of that, they had to rate several variables on seven-point Likert scales to measure for confounding factors. These included the AI usefulness, their experience with AI, their cognitive load, and their trust in AI. All variables are presented in Appendix A in Table 5. Throughout the questionnaire, we implemented a second attention check to ensure only valid results, as suggested by Abbey and Meloy (2017). In this attention check, we asked participants to select ”Likely“ for an item of AI usability (”As an attention check, please choose ”Likely” for this statement“). The questionnaire ended with demographic questions on participants’ age, gender, education and employment.
@@ 4.3. AI Development
The AI used in the study was a LLM that provided participants with a classification (AI advice) and an explanation for its reasoning, depending on the treatment they were assigned to. As LLM, we used Open AI’s GPT-4o model (model version 2024-05-13) through an Azure Open AI Studio instance. In the pre-phase of the study, we tested multiple prompt strategies in a workshop with three authors. The final prompts that were used are shown in Appendix A in Table 4. The LLM was prompted such that the correct explanation treatment provided an explanation that corresponded to its prediction, while the explanation did not match the prediction in the incorrect explanation treatment.
The definition of an incorrect explanation follows Morrison et al. (2024) and Cabitza et al. (2024a) in that the AI does not provide correct reasoning for its predicted class (not coherent with the AI advice), independent of the ground truth of the image. For instance, if the AI made a correct prediction for an Art Nouveau building, but the explanation does not conform to this class, it is defined as incorrect. The definition further follows Cabitza et al. (2024a) in that the incorrect explanations are pertinent as they emphasize features that are not relevant for a given predicted architecture style. For instance, if the incorrect explanation highlights features of an Art Nouveau building that are actually characteristic of Art Deco. To avoid information overload (Arnold et al., 2023), we designed the explanations to be no longer than three sentences. As outlined in Section 3, we analyze the impact of incorrect explanations for cases where the AI classification is correct.
@@ 4.5. Metrics
Similar to previous work (e.g., (Hemmer et al., 2023; Schoeffer et al., 2024)), we assessed participants’ task performance in classifying the architectural styles in pre-test, main task, and post-test. Aligning with prior research (McCormick, 1997; Nahdi and Jatisunda, 2020), we use the task performance to infer participants’ procedural knowledge. We measure the human-AI team performance in the main task. As metrics for the task performance, we used accuracy and measured the ratio of correctly classified images over all images. Participants had to select one of the three different architectural styles for each image by selecting from a drop-down menu on each page of the task. This means that a random guess corresponded to 33.3% of performance. Aligning with Schoeffer et al. (2024), we also measured the correct adherence and detrimental overrides in the main task of the study.
We assessed participants’ reasoning ability for the three architectural styles through open-ended questions by asking them to describe and explain each style, thereby following the procedure of Chi et al. (1989). We assessed participants’ reasoning abilities in terms of correctness by comparing them to the correct characteristics and features for each architectural style. We followed the procedure of Huang et al. (2024) and used LLMs to evaluate the correctness of participants’ answers and computed reasoning scores for participants: zero represents a completely incorrect or irrelevant answer, and five represents a completely correct and comprehensive answer. The prompt used for this analysis is in Table 4 in Appendix A. By doing so, we were able to assess whether incorrect explanations impaired participants’ reasoning.
Finally, we established several control variables to investigate the potential underlying factors that might influence AI-assisted decision-making. In particular, we controlled for participants’ cognitive load as previous research suggests that the information in explanations displayed to humans can affect their decision-making behavior (Abdul et al., 2020; Hudon et al., 2021; Herm, 2023; Spitzer et al., 2024a). We measured participants’ cognitive load on a seven-point Likert scale by having them rate five validated items previously developed by Hart (1986) and that were already applied in the field (Senoner et al., 2024). In addition, we assessed participants’ AI trust, AI usefulness, and experience with AI by using items proven in previous research (Senoner et al., 2024), all of them also rated on a seven-point Likert scale. All items are shown in Table 5 in Appendix A.
@@ 5. Results
It took participants, on average, 14 minutes and 54 seconds to complete the study. Overall, 160 participants passed the attention check and finished the study according to the study protocol. Of these 160 participants, 79 were male, 77 were female, and four identified as diverse. To address the research questions, we first conduct several statistical analyses to answer RQ: 1 in subsection 5.1. Subsequently, we address RQ: 2 and qualitatively assess the open-ended questionnaires in subsection 5.2. In the final subsection 5.3, we evaluate the impact of incorrect explanations on the human-AI team performance (RQ: 3).
@@ 5.1. RQ1: Impact of Incorrect Explanations on Procedural Knowledge
To establish a baseline and ensure that all treatments began on an equal performance level in classifying the architectural styles, we conduct a one-way ANOVA on the pre-test performance scores across the four treatments (). These results fail to reject the null hypothesis, indicating no significant differences in pre-test performance among the four treatments. This finding suggests that participants across all treatments start with comparable levels of procedural knowledge to classify the architectural styles prior to the main task. This allows for a thorough interpretation of any differences observed in the post-test results, as they can be more readily attributed to the treatment effects rather than pre-existing differences.
@@
To assess the significance of differences in post-test performance across treatments, we compare performance levels between treatments. A one-way ANOVA yields evidence of treatment effects (), indicating that the type of AI support in the main task influences participant performance in the post-test. To further examine these differences, we conduct pairwise comparisons using one-sided t-tests, assuming participants in the incorrect explanation treatment exhibit lower performance compared to participants in the other treatment. We report the corrected p-values according to the Holm-Bonferroni correction (Holm, 1979) (see Table 8 in Appendix A). There is no significant difference between the post-test performances in the incorrect explanation treatment and the control group (). Comparison with the AI treatment also shows no significant difference (). Notably, the comparison between incorrect and correct explanations yields a significant difference (), with incorrect explanations leading to inferior performance. With these findings, we find support for hypothesis 1. These results demonstrate that the correctness of explanations plays an important role in AI-assisted decision-making, with incorrect explanations decreasing the task performance of humans in the post-test the most. Correct explanations provide a performance advantage over incorrect explanations, underscoring the importance of explanation accuracy in AI-assisted tasks. This is an interesting finding: we compare the post-test performance levels between the treatments in a one-sided manner and assume that the incorrect explanation of treatment leads to a performance level lower than that of the other treatments. While we can see that participants provided with incorrect explanations perform worse, correct explanations lead to the highest post-test performance. Contrarily, compared to having no explanation or no AI advice at all, this effect diminishes, potentially showing that incorrect explanations do not lower the post-test performance as much as when correct explanations are provided.
@@
To identify the potential reasons behind these findings, we further analyze which control variables influence the post-test performance. We run a regression analysis with the post-test performance as the dependent variable and model the type of AI support as the independent variable. We define cognitive load, AI trust, AI knowledge, and AI usefulness as confounding factors. The regression analysis examining the impact of different treatments on post-test performance reveals that the type of AI support significantly influences participants’ performance (see Table 9 in Appendix A). The model accounts for of the variance in post-test performance (). Specifically, participants in the correct explanation treatment exhibit the greatest improvement in post-test performance, with a significant positive effect (), followed by those in the AI treatment () and the incorrect explanation treatment (). The data supports the findings of the statistical tests. Interestingly, cognitive load is negatively associated with post-test performance (), suggesting that higher cognitive load impedes procedural knowledge. We further analyze this finding by comparing the cognitive load for each treatment. We can see that cognitive load is highest for the control group () and the incorrect explanation treatment () compared to the other treatments (AI treatment = , correct explanations treatment = ) illustrating that the decrease in procedural knowledge stems from the increased cognitive load. Contrarily, trust, AI knowledge, and AI usefulness do not impact the post-test performance.
@@
Additionally, we analyze how the procedural knowledge develops from pre-test to post-test. We plot this difference in Figure 4. The figure illustrates whether procedural knowledge develops or degrades across different decision-making stages in the different treatments. Interestingly, the data reveals that the incorrect explanation treatment still leads to procedural knowledge development. Conversely, the AI treatment and correct explanation treatment show the highest procedural knowledge development, exceeding the development in the incorrect explanation treatment.
@@
Furthermore, we conduct an ANOVA and Tukey HSD Post-Hoc Test which reveal significant differences in procedural knowledge gains across treatments (). Notably, the incorrect explanation treatment does not significantly differ from the AI treatment or the correct explanation treatment ( for both comparisons), but it shows a trend toward higher procedural knowledge gains compared to the control group (). The procedural knowledge development is significantly higher than in the control group (). This suggests that while incorrect explanations may increase procedural knowledge in collaboration with an AI, the correctness of explanations reduces the effect compared to correct explanations.